OpenStudy (shiree19):

Given the function rule f(x) = x^2 - 3x + 2, what is the output of f(-2)? Choices: A - -8 B - 0 C - 4 D - 12 I selected B - 0. That is wrong. The correct answer is D - 12. How do you get this?

OpenStudy (solomonzelman):

f(x) = x^2 - 3x + 2 f(4) = (4)^2 - 3(4) + 2= 16-12+2 = 6 f(-2) = ?

OpenStudy (solomonzelman):

plug in -2, just like I plugged in 4

OpenStudy (shiree19):

ok. just a sec.

OpenStudy (solomonzelman):

sure

OpenStudy (shiree19):

Am I right? f(-2) = -2^2 = 4 4 - (3 * -2) = 4 - (-6) = 10 + 2 = 4

OpenStudy (shiree19):

I meant 10 + 2 = 12 sorry

OpenStudy (solomonzelman):

yeah, 12 is right

OpenStudy (shiree19):

thanks so much!

OpenStudy (solomonzelman):

My pleasure !